a motorcycle accelerates from rest for 6.0 seconds at 4.0 m/s [forwards] and then the acceleration stops and it travels at a constant velocity for another 4.0 seconds. what is the velocity of the motorcycle at the end of the first 6.0 seconds? How far does the motorcycle travel in the 10 seconds?
1
Expert's answer
2019-10-25T11:01:58-0400

The motorcycle at the end of the first 6.0 seconds:


"v=at=(4)(6)=24\frac{m}{s}"

The distance will be:


"s=0.5at^2+vT=0.5(4)(6)^2+(24)4=170\ m"